# KEPLA: A Knowledge-Enhanced Deep Learning Framework for Accurate Protein-Ligand Binding Affinity Prediction

### v1

@@ Results
For complex 3UI7 (pyrazoloquinoline bound to human PDE10A), the ligand molecule pyrazoloquinoline features a fused structure composed of a pyrazole ring and a quinoline ring. The pyrazole ring is a five-membered nitrogen-containing heterocycle, while the quinoline ring consists of a pyridine ring fused with a benzene ring. Our model correctly identifies the nitrogen-containing heterocyclic region as critical for protein-ligand binding: the nitrogen atom bearing a hydrogen in the pyrazole ring acts as a hydrogen bond donor to the side chain of Gln729; the other nitrogen atom interacts with Thr685 via a solvent molecule; and the nitrogen atom in the pyridine ring serves as a hydrogen bond acceptor to Gln729. Conversely, the methoxy group on the benzene ring is incorrectly predicted to form specific interactions with the protein, despite being solvent-exposed, though this exposure may facilitate additional binding interactions (blue highlight). Additionally, the pyridine ring, which engages in an arene-arene interaction with Phe729, is only partially emphasized by the interpretability model. Notably, although only five atoms of pyrazoloquinoline are among the top 20% atoms, all correspond to distinct protein-ligand interaction sites confirmed by the X-ray crystal structure.
In complex 2WTV (MLN8054 bound to human Aurora kinase A), the core scaffold of MLN8054 comprises a fused polycyclic structure connected to a difluorophenyl group and an anilino moiety, with chlorine and carboxylate substituents. The interpretability feature of our model effectively highlights key interaction patterns relevant to protein-ligand binding. For instance, within the aniline group, the amino group is correctly identified as forming a hydrogen bond with the backbone of Ala213 (acting as a donor), and the adjacent benzene ring is partially highlighted due to its arene-hydrogen interaction with Gly216. The carboxylate anion is also accurately identified, interacting with Tyr212 via a solvent-mediated contact. Furthermore, the nitrogen-containing aromatic heterocycle is recognized for its potential to act as either a hydrogen bond donor or acceptor, with one specific nitrogen atom explicitly highlighted for donating a hydrogen bond to the backbone of Glu211. The fluorine atoms in the difluorophenyl group and the chlorine substituent are also marked, likely due to their solvent exposure and potential role in enhancing binding affinity. Despite MLN8054 (in 2WTV) being considerably larger and more complex than pyrazoloquinoline (in 3UI7), the model demonstrates strong interpretability by accurately reflecting most of the key interactions confirmed by the crystal structure.
@@
Regarding the more challenging task of interpreting protein sequences, the results are less specific compared to those for ligands. This limitation arises from the model’s design, where attention weights are computed and visualizations are generated based on substructural units composed of multiple residues. While this design choice improves computational efficiency, it inevitably reduces interpretability at the residue level, as it does not explicitly resolve which specific residues within each substructure are responsible for binding. Nonetheless, several key amino acid residues involved in binding are correctly highlighted (Figure 4b). For example, in complex 3UI7, residues Gln726 and Phe729, both forming specific interactions with the ligand, are accurately highlighted, along with additional pocket residues such as Tyr524, His525, Leu675, Ser677, and Val678. In complex 2WTV, the model correctly highlights Glu211 and Ala213, which participate directly in ligand binding, as well as several additional binding-site residues: Leu139, Gly140, Lys141, Gly142, Ala160, Leu194, Leu210, and Tyr212. In complex 2QE4, Glu353 and His524, key hydrogen bond partners of the ligand, are successfully highlighted, in addition to pocket-forming residues Thr347, Leu349, Ala350, Met421, Ile424, Gly521, and Leu525. Statistically, across the three complexes, nearly half of the binding-site residues are correctly identified by the model. These results suggest that substructure-level attention provides a promising foundation for interpretability. Incorporating residue-level attention mechanisms could further enhance the resolution and accuracy of protein-ligand interaction explanations. Moreover, if predictions are not constrained to be interaction-free, integrating 3D structural and local spatial information from the binding pocket may further improve interpretability.
@@
Ablation studies have confirmed that incorporating external knowledge related to proteins and ligands significantly improves affinity prediction. Building on this, we further investigate whether such external knowledge can also enhance the interpretability of protein-ligand binding. To this end, we identify the entities most closely associated with the target protein and ligand and attempt to derive plausible explanations for their interaction based on these entities. Specifically, we utilize the learned embeddings of heads, relations, and tails from the KG, and continue applying the scoring function to evaluate all triples involving the target protein or ligand. A lower score indicates greater semantic proximity between the protein or ligand and the corresponding entity. As shown in Figure 5, the top five closest entities are presented. These entities are considered the most relevant and are used, along with their definitions, to interpret the protein-ligand binding mechanism. To maintain consistency with prior experiments, we use the same case studies: PDB IDs 3UI7, 2WTV, and 2QE4.
@@
For the PDB structure 2WTV (MLN8054 bound to human Aurora kinase A), the top GO entities most closely associated with the protein include protein kinase binding, cytosol, protein serine kinase activity, ATP binding, and protein phosphorylation. These annotations indicate that Aurora kinase A plays a crucial role in spindle assembly and chromosome segregation during mitosis, primarily through its ATP-binding and serine/threonine kinase activities within the cytoplasm. The interaction between the ligand MLN8054 and Aurora kinase A is highly consistent with these functional annotations. MLN8054 is a selective inhibitor of Aurora kinase A that mimics ATP and competitively binds to its ATP-binding site. This binding inhibits the kinase’s activity, suppressing phosphorylation of downstream substrate proteins and thereby disrupting cell division. On the ligand side, the entities most closely associated with MLN8054 include acidic group, three aromatic carbocycles, four rotatable bonds, nine heteroatoms, and one aromatic heterocycle. As shown in Figure 4a, the presence of heteroatoms facilitates hydrogen bonding and potential metal coordination, contributing to specific and high-affinity interactions with the protein. Aromatic heterocycles, commonly found in high-affinity drugs, are especially effective in optimizing molecular compatibility with binding pockets. Moreover, a greater number of rotatable bonds enhances the ligand’s conformational flexibility, allowing it to better fit the shape of the binding site. Together, these molecular features help explain the strong binding affinity observed in the 2WTV complex.
@@
For the PDB structure 2QE4 (benzopyran agonist complexed with the human estrogen receptor), the top GO entities most closely associated with the protein are nuclear receptor activity, steroid binding, positive regulation of transcription by RNA polymerase II, estrogen response element binding, and nucleus. Notably, steroid binding highlights the role of the estrogen receptor as a steroid hormone receptor that can be activated by benzopyran agonists exhibiting steroid-like structural features. The remaining GO annotations indicate that, upon activation, the estrogen receptor translocates into the nucleus, binds to specific DNA sequences known as estrogen response elements, and regulates the transcription of target genes involved in processes such as cell proliferation and differentiation. On the ligand side, the most relevant molecular entities for the benzopyran agonist in 2QE4 include two hydrogen bond donors, one aliphatic heterocycle, three rotatable bonds, two aromatic carbocycles, and hydrophobe. Among these, the feature of two hydrogen bond donors is consistent with the interaction profile shown in Figure 4a. Hydrophobicity is also a critical factor in protein-ligand binding, as it promotes favorable interactions between nonpolar groups, contributing to the stabilization of the binding site and enhancing overall binding affinity. The implications of the remaining features have been discussed in the earlier examples and similarly contribute to binding specificity and adaptability.
@@
The above examples demonstrate that the KEPLA model can effectively retrieve knowledge entity nodes associated with both the protein and the ligand, thereby leveraging biologically and chemically grounded textual knowledge to explain their binding interactions. Compared to purely data-driven black-box predictions, this knowledge-based interpretability provides a more tangible and trustworthy foundation for rational drug discovery.
@@ Discussion
In this work, we present KEPLA, an end-to-end knowledge-enhanced deep learning framework for protein-ligand binding affinity (PLA) prediction, designed to address the performance bottlenecks and limited interpretability of existing methods. Specifically, we construct biochemical KGs for proteins and ligands, respectively. The encoded representations of proteins and ligands are integrated into KG embedding objectives to capture the rich semantic information within these graphs. Subsequently, the model employs cross attention mechanisms to perform fine-grained aggregation of the knowledge-enhanced representations, which are then fed into a decoder to predict binding affinity. We conduct extensive experiments to evaluate the effectiveness of KEPLA in addressing the target challenges.
@@
To assess predictive performance, we benchmark KEPLA against several state-of-the-art baselines (spanning both interaction-free and interaction-based methods) on two widely used public datasets. Our experimental setup includes both a random split (to test in-domain performance) and a cluster-based split (to assess cross-domain generalization). Compared to state-of-the-art baselines, KEPLA consistently achieves superior PLA prediction accuracy under both settings. Notably, despite being an interaction-free model, KEPLA outperforms many interaction-based methods in the in-domain setting and maintains robust performance in the more challenging cross-domain scenario. Ablation studies further confirm the critical role of KGs in enhancing model performance. These findings suggest that incorporating external domain knowledge not only improves prediction accuracy but also mitigates the hidden data bias inherent in purely data-driven approaches.
@@
To evaluate the interpretability of our model, we conduct case studies that leverage both the cross attention and KG modules to provide explanations from structural and knowledge-based perspectives. By mapping attention weights to protein subsequences and ligand atoms, KEPLA can highlight interacting regions such as the protein binding pocket and ligand atoms, offering structural insights into the predicted interactions. Furthermore, by linking proteins and ligands to GO and LP entities in the KG, our model grounds predictions in biochemical context, thereby enhancing interpretability from a knowledge-driven perspective.
@@
This work focuses on interaction-free PLA prediction using 1D protein sequences and 2D molecular graphs as input. Given that high-quality 3D protein structures are available for only a limited subset of known sequences, we deliberately refrain from incorporating 3D structural modeling in this study. However, recent breakthroughs, such as AlphaFold by DeepMind [46, 47], which has predicted over 2 billion protein structures across 1 million species, present compelling opportunities for integrating 3D structural information into future chemogenomics-based PLA models. Inspired by principles of pairwise local interaction learning and domain adaptation, we anticipate that extending our approach to include 3D structures could further enhance both predictive performance and model interpretability. Furthermore, our current KG construction is grounded in basic biochemical knowledge. Expanding this to incorporate richer domain-specific and cross-disciplinary knowledge offers a promising direction for future improvement. Finally, the KEPLA framework is highly generalizable and can be readily adapted to other molecular interaction prediction tasks, including drug-drug interaction [48] and protein-protein interaction [49].
@@ Methods
KEPLA consists of four main components: (1) ESM-based protein sequence encoding for learning protein representations; (2) GCN-based molecular graph encoding for learning ligand representations; (3) KG embedding objective, enhancing representations through external biochemical knowledge graph (KG); (4) PLA prediction objective, enabling fine-grained modeling of protein-ligand interactions via cross attention.
Problem formulation
In PLA prediction, the goal is to estimate the binding strength between a target protein and a ligand compound. We represent each protein sequence as $\mathcal{P}=(a_{1},a_{2},...,a_{n})$, where each $a_{i}$ denotes one of the 23 amino acids. For the ligand compound, most existing deep learning approaches use the SMILES [7], a 1D string format that encodes atoms and bond information in a molecule. While SMILES facilitates the use of classical deep learning architectures, its linear format does not naturally capture molecular structures, potentially leading to the loss of crucial spatial information and reduced predictive accuracy. To address this, our model converts each SMILES string into a corresponding 2D molecular graph. Formally, a ligand is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes (atoms), and $\mathcal{E}$ is the set of edges (chemical bonds) [50]. Given a protein sequence $\mathcal{P}$ and a ligand graph $\mathcal{G}$, the PLA prediction task aims to learn a model $\mathcal{M}$ that maps the joint representation space $\mathcal{P}\times\mathcal{G}$ to a continuous binding affinity value $y\in\mathbb{R}$. The commonly used notations in this paper are summarized in the Supplementary Information.
@@
The protein encoder is built upon the pre-trained ESM model [18], augmented with a multi-layer neural network. It transforms an input amino acid sequence into a set of latent vectors, each representing a local subsequence of the protein. ESM, a widely recognized protein language model, has demonstrated strong performance in predicting protein structure, function, and other properties from individual sequences. To fully leverage ESM’s powerful encoding capabilities, we use it as the backbone of our protein encoder. Given a target protein sequence $\mathcal{P}$, we input it into the pre-trained ESM model and extract the final hidden state as the initial feature matrix:
@@
$\mathbf{X}_{p}=\text{AvgPool}_{s}(\mathbf{M}_{p}).$ (2)
Next, a multi-layer feedforward neural network is employed to adapt these features to our task and extract higher-level abstractions of local protein fragments. The encoding process is defined as:
@@
where $\mathbf{W}_{n}^{(l)}$ and $\mathbf{b}_{n}^{(l)}$ are the learnable weights and biases of the $l$-th layer, $\mathbf{H}_{p}^{(l)}$ is the $l$-th hidden protein representation, and $\sigma(\cdot)$ is a non-linear activation function (ReLU in our case). The initial input is set as $\mathbf{H}_{p}^{(0)}=\mathbf{X}_{p}$. The final output, $\mathbf{H}_{p}\in\mathbb{R}^{D\times M}$, is adopted as the protein representation, where $D$ is the representation dimension and $M$ is the number of local fragments.
@@
For each ligand compound, we convert its SMILES string into a 2D molecular graph $\mathcal{G}$. To represent the node (atom) information in $\mathcal{G}$, we initialize each atom using its chemical properties, following the implementation in the DGL-LifeSci library [51]. Each atom is initially represented as a 74-dimensional integer vector describing eight pieces of information: atom type, degree, number of implicit hydrogens, formal charge, number of radical electrons, hybridization state, number of total hydrogen atoms, and aromaticity:
@@
This yields the initial node feature matrix $\mathbf{M}_{d}\in\mathbb{R}^{74\times N}$, where $N$ is the maximum allowed number of atoms per molecule. Molecules with fewer atoms are padded with virtual nodes filled with zeros to ensure consistent dimensions for batch training. Next, a linear transformation is applied to project the integer feature matrix into a real-valued dense representation:
@@
producing an input feature matrix $\mathbf{X}_{d}\in\mathbb{R}^{D_{d}\times N}$, with $D_{d}$ denoting the output dimension.
@@
where $\mathbf{W}_{g}^{(l)}$ and $\mathbf{b}_{g}^{(l)}$ are the learnable weight matrix and bias vector of the $l$-th GCN layer, $\mathbf{H}_{d}^{(l)}$ is the $l$-th hidden node representation, $\mathbf{A}$ is the adjacency matrix of the graph $\mathcal{G}$ (augmented with self-loops), and $\sigma(\cdot)$ denotes a non-linear activation function (ReLU in our case). The input to the first layer is set as $\mathbf{H}_{d}^{(0)}=\mathbf{X}_{d}$. The final output, $\mathbf{H}_{d}\in\mathbb{R}^{D\times N}$, is used as the ligand representation.
@@
Thus far, the two encoders described above only capture the 1D and 2D structural information of proteins and ligands, respectively. To further enrich these representations with biochemical factual knowledge, we integrate them into a KG embedding framework. Specifically, we treat the encoded protein and ligand representations as entity embeddings within a KG and incorporate them into a joint learning process alongside other entities and their relations. The remainder of this subsection details our implementation.
Our KG structurally consists of entity-relation-entity triples to describe relational knowledge. We define a triplet as $(h,r,t)$, where $h$ and $t$ are head and tail entities, and $r$ is the relation linking them, whose type is usually predefined by the schema. We construct two distinct subgraphs: one centered around proteins and the other around ligands.
  • Protein KG (GO-based): This subgraph incorporates two types of entities: protein entities $e_{protein}$ involved in the affinity prediction task, and Gene Ontology (GO) entities $e_{GO}$ describing functional annotations. The GO entities cover three categories: Molecular Function (MF), Cellular Component (CC), and Biological Process (BP). Therefore, the protein-related triples fall into three groups: protein-MF, protein-CC, and protein-BP.
  • Ligand KG (LP-based): Similarly, this subgraph includes ligand entities $e_{ligand}$ and ligand property (LP) entities $e_{LP}$. The LP entities are further divided into Molecular Descriptors (MD), i.e., numerical attributes derived from molecular structures (e.g., counts of hydrogen bond donors/acceptors), and Chemical Features (CF), which denote qualitative attributes (e.g., hydrophobe, lumped-hydrophobe). Accordingly, the ligand-related triples fall into two groups: ligand-MD and ligand-CF.
Next, we involve the protein and ligand representations in the KG embedding objective to enhance the protein and ligand encoding with external knowledge. Since the encoders yield local matrix representations unsuitable for KG embeddings, we first compute global vector representations by averaging across the local feature dimensions, then project them into the KG embedding space via a linear transformation:
@@
where $\text{Average}(\cdot)$ denotes the averaging operation over all local fragment representations. $\mathbf{W}_{k}^{*}$ and $\mathbf{b}_{k}^{*}$ are learnable weight matrices and bias vectors that map the global representations into the KG embedding space. The resulting embeddings $\mathbf{h}_{p}$ and $\mathbf{h}_{d}\in\mathbb{R}^{D}$ serve as the semantic representations of the protein and ligand entities in the KG. Similar to [52], the embedding sets of other entities and the relations are denoted by learnable embeddings $\mathbf{T}\in\mathbb{R}^{D\times T}$ and $\mathbf{R}\in\mathbb{R}^{D\times R}$, where $T$ is the number of non-protein and non-ligand entities, and $R$ is the number of relation types.
To achieve the KG embedding objective, we train the KG by optimizing the embeddings to fit the triplet relationships. The objective function is defined as:
$\mathcal{L}_{KE}=\frac{1}{|\mathcal{S}|}\sum_{(h,r,t)\in\mathcal{S}}F(h,r,t),$ (8)

### v2

@@ Results
For complex 3UI7 (pyrazoloquinoline bound to human PDE10A), the ligand molecule pyrazoloquinoline features a fused structure composed of a pyrazole ring and a quinoline ring. The pyrazole ring is a five-membered nitrogen-containing heterocycle, while the quinoline ring consists of a pyridine ring fused with a benzene ring. Our model correctly identifies the nitrogen-containing heterocyclic region as critical for protein-ligand binding: the nitrogen atom bearing a hydrogen in the pyrazole ring acts as a hydrogen bond donor to the side chain of Gln729; the other nitrogen atom interacts with Thr685 via a solvent molecule; and the nitrogen atom in the pyridine ring serves as a hydrogen bond acceptor to Gln729. Conversely, the methoxy group on the benzene ring is incorrectly predicted to form specific interactions with the protein, despite being solvent-exposed, though this exposure may facilitate additional binding interactions (blue highlight). Additionally, the pyridine ring, which engages in an arene-arene interaction with Phe729, is partially emphasized by the interpretability model. Notably, although only five atoms of the ligand are among the top 20% atoms, most correspond to distinct protein-ligand interaction sites confirmed by the X-ray crystal structure.
In complex 2WTV (MLN8054 bound to human Aurora kinase A), the core scaffold of MLN8054 comprises a fused polycyclic structure connected to a difluorophenyl group and an anilino moiety, with chlorine and carboxylate substituents. The interpretability of our model effectively highlights key interaction patterns relevant to protein-ligand binding. For instance, within the aniline group, the amino group is correctly identified as forming a hydrogen bond with the backbone of Ala213 (acting as a donor), and the adjacent benzene ring is partially highlighted due to its arene-hydrogen interaction with Gly216. The carboxylate anion is also accurately identified, interacting with Tyr212 via a solvent-mediated contact. Furthermore, the nitrogen-containing aromatic heterocycle is recognized for its potential to act as either a hydrogen bond donor or acceptor, with one specific nitrogen atom explicitly highlighted for donating a hydrogen bond to the backbone of Glu211. The fluorine atoms in the difluorophenyl group and the chlorine substituent are also marked, likely due to their solvent exposure and potential role in enhancing binding affinity. Despite MLN8054 (in 2WTV) being considerably larger and more complex than pyrazoloquinoline (in 3UI7), the model demonstrates strong interpretability by accurately reflecting most of the key interactions confirmed by the crystal structure.
@@
Regarding the more challenging task of interpreting protein sequences, the results are less specific compared to those for ligands. This limitation arises from the model’s design, where attention weights are computed and visualizations are generated based on protein fragments composed of multiple residues. While this design choice improves computational efficiency, it inevitably reduces interpretability at the residue level, as it does not explicitly resolve which specific residues within each fragment are responsible for binding. Nonetheless, several key amino acid residues involved in binding are correctly highlighted as shown in Figure 4b. For example, in complex 3UI7, residues Gln726 and Phe729, both forming specific interactions with the ligand, are accurately highlighted, along with additional pocket residues such as Tyr524, His525, Leu675, Ser677, and Val678. In complex 2WTV, our model correctly highlights Glu211 and Ala213, which participate directly in ligand binding, as well as several additional binding-site residues: Leu139, Gly140, Lys141, Gly142, Ala160, Leu194, Leu210, and Tyr212. In complex 2QE4, Glu353 and His524 are successfully highlighted as key hydrogen bond partners of the ligand, in addition to pocket-forming residues Thr347, Leu349, Ala350, Met421, Ile424, Gly521, and Leu525. Statistically, across the three complexes, nearly half of the binding-site residues are correctly identified by the model. These results suggest that fragment-level attention provides a promising foundation for interpretability. Incorporating residue-level attention mechanisms could further enhance the resolution and accuracy of protein-ligand interaction explanations. Moreover, if predictions are not constrained to be interaction-free, integrating 3D structural and local spatial information from the binding pocket may further improve interpretability.
@@
Ablation studies have confirmed that incorporating external knowledge related to proteins and ligands significantly improves affinity prediction. Building on this, we further investigate whether such external knowledge can also enhance the interpretability of protein-ligand binding. To this end, we identify the entities most closely associated with the target protein and ligand and attempt to derive plausible explanations for their interaction based on these entities. Specifically, we utilize the learned embeddings of entities and relations from the KG and continue applying the scoring function to evaluate all triplets involving the target protein or ligand. A lower score indicates greater semantic proximity between the protein or ligand and the corresponding GO or LP entity. As shown in Figure 5, the top five closest entities are presented. These entities are considered the most relevant and are used, along with their definitions, to interpret the protein-ligand binding mechanism. To maintain consistency with prior experiments, we use the same case studies: PDB IDs 3UI7, 2WTV, and 2QE4.
@@
For the PDB structure 2WTV (MLN8054 bound to human Aurora kinase A), the top GO entities most closely associated with the protein include protein kinase binding, cytosol, protein serine kinase activity, ATP binding, and protein phosphorylation. These annotations indicate that Aurora kinase A plays a crucial role in spindle assembly and chromosome segregation during mitosis, primarily through its ATP-binding and serine/threonine kinase activities within the cytoplasm. The interaction between the ligand MLN8054 and Aurora kinase A is highly consistent with these functional annotations. MLN8054 is a selective inhibitor of Aurora kinase A that mimics ATP and competitively binds to its ATP-binding site. This binding inhibits the kinase’s activity, suppressing phosphorylation of downstream substrate proteins and thereby disrupting cell division. On the ligand side, the LP entities most closely associated with MLN8054 include acidic group, three aromatic carbocycles, four rotatable bonds, nine heteroatoms, and one aromatic heterocycle. As shown in Figure 4a, the presence of heteroatoms facilitates hydrogen bonding and potential metal coordination, contributing to specific and high-affinity interactions with the protein. Aromatic heterocycles, commonly found in high-affinity drugs, are especially effective in optimizing molecular compatibility with binding pockets. Moreover, a greater number of rotatable bonds enhances the ligand’s conformational flexibility, allowing it to better fit the shape of the binding site. Together, these molecular features help explain the strong binding affinity observed in the 2WTV complex.
@@
For the PDB structure 2QE4 (benzopyran agonist complexed with the human estrogen receptor), the top GO entities most closely associated with the protein are nuclear receptor activity, steroid binding, positive regulation of transcription by RNA polymerase II, estrogen response element binding, and nucleus. Notably, steroid binding highlights the role of the estrogen receptor as a steroid hormone receptor that can be activated by benzopyran agonists exhibiting steroid-like structural features. The remaining GO annotations indicate that, upon activation, the estrogen receptor translocates into the nucleus, binds to specific DNA sequences known as estrogen response elements, and regulates the transcription of target genes involved in processes such as cell proliferation and differentiation. On the ligand side, the most relevant LP entities for the benzopyran agonist in 2QE4 include two hydrogen bond donors, one aliphatic heterocycle, three rotatable bonds, two aromatic carbocycles, and hydrophobe. Among these, the feature of two hydrogen bond donors is consistent with the interaction profile shown in Figure 4a. Hydrophobicity is also a critical factor in protein-ligand binding, as it promotes favorable interactions between nonpolar groups, contributing to the stabilization of the binding site and enhancing overall binding affinity. The implications of the remaining features have been discussed in the earlier examples and similarly contribute to binding specificity and adaptability.
@@
The above examples demonstrate that the KEPLA model can effectively retrieve knowledge entities associated with both the protein and the ligand through KG, thereby leveraging biologically and chemically grounded textual knowledge to explain their binding interactions. Compared to purely data-driven black-box predictions, this knowledge-based interpretability can provide a more tangible and trustworthy foundation for rational drug discovery.
@@ Discussion
In this work, we present KEPLA, an end-to-end knowledge-enhanced deep learning framework for protein-ligand binding affinity (PLA) prediction, designed to address the performance bottlenecks and limited interpretability of existing methods. Specifically, we construct a biochemical KG for proteins and ligands, respectively. The encoded representations of proteins and ligands are integrated into the KG embedding objective to capture the rich semantic information within KG. Subsequently, the model employs cross attention mechanisms to perform fine-grained aggregation of the knowledge-enhanced representations, which are then fed into a decoder to predict binding affinity. We conduct extensive experiments to evaluate the effectiveness of KEPLA in addressing the performance and interpretability challenges.
@@
To assess the predictive performance of our model, we compare KEPLA against several state-of-the-art baselines (spanning both interaction-free and interaction-based methods) on two widely used public datasets. Our experimental setup includes both a random split to assess in-domain performance and a cluster-based split to assess cross-domain generalization. Compared to state-of-the-art baselines, KEPLA consistently achieves superior PLA prediction accuracy under both settings. Notably, despite being an interaction-free model, KEPLA outperforms many interaction-based methods in the in-domain setting and maintains robust performance in the more challenging cross-domain scenario. Ablation study further confirms the critical role of KG in enhancing model performance. These findings suggest that incorporating external domain knowledge not only improves prediction accuracy but also mitigates the hidden data bias inherent in purely data-driven approaches.
@@
To evaluate the interpretability of our model, we conduct case studies that leverage both the cross attention and KG modules to provide explanations from structural and knowledge-based perspectives. By mapping attention weights to protein fragments and ligand atoms, KEPLA can highlight interacting regions such as the protein binding pocket and ligand interacting atoms, offering structural insights into the predicted interactions. Furthermore, by linking proteins and ligands to GO and LP entities in the KG, KEPLA grounds predictions in biochemical context, thereby enhancing interpretability from a knowledge-driven perspective.
@@
This work focuses on interaction-free PLA prediction using 1D protein sequences and 2D molecular graphs as input. Given that high-quality 3D protein structures are available for only a limited subset of known sequences, we deliberately refrain from incorporating 3D structural information in this study. However, recent breakthroughs, such as AlphaFold by DeepMind [48, 49], which has predicted over 2 billion protein structures across a million species, present compelling opportunities for integrating 3D structural information into future chemogenomics-based PLA models. Inspired by principles of pairwise local interaction learning and domain adaptation, we anticipate that extending our approach to include 3D structures could further enhance both predictive performance and model interpretability. Furthermore, our current KG construction is grounded in basic biochemical knowledge. Expanding this to incorporate richer domain-specific and cross-disciplinary knowledge offers a promising direction for future improvement. Finally, the KEPLA framework is highly generalizable and can be readily adapted to other molecular interaction prediction tasks, including drug-drug interaction [50] and protein-protein interaction [51].
@@ Methods
Problem formulation
In PLA prediction, the goal is to estimate the binding strength between a target protein and a ligand compound. We represent each protein sequence as $\mathcal{P}=(a_{1},a_{2},...,a_{K})$, where each $a_{i}$ denotes one of the 23 amino acids. For the ligand compound, most existing deep learning approaches use the SMILES [9], a 1D string format that encodes atoms and bond information in a molecule. While SMILES facilitates the use of classical deep learning approaches, its linear format does not naturally capture molecular structures, potentially leading to the loss of crucial spatial information and reduced predictive accuracy. To address this, our model converts each SMILES string into a corresponding 2D molecular graph. Formally, a ligand is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes (atoms), and $\mathcal{E}$ is the set of edges (chemical bonds) [52]. Given a protein sequence $\mathcal{P}$ and a ligand graph $\mathcal{G}$, the PLA prediction task aims to learn a model $\mathcal{M}$ that maps the joint representation space $\mathcal{P}\times\mathcal{G}$ to a continuous binding affinity value $y\in\mathbb{R}$. The commonly used notations in this paper are summarized in the Supplementary Information.
@@
As a PLA prediction model, KEPLA consists of four main components: (1) ESM-based protein sequence encoding for learning protein representations; (2) GCN-based molecular graph encoding for learning ligand representations; (3) KG embedding objective, enhancing representations through external biochemical KG; (4) PLA prediction objective, enabling fine-grained modeling of protein-ligand interactions via cross attention. Subsequently, we present a detailed elucidation of these components.
@@
The protein encoder is built upon the pre-trained ESM model [20], augmented with a multi-layer neural network. It transforms an input amino acid sequence into a set of latent vectors, each representing a local fragment of the protein. ESM, a widely recognized protein language model, has demonstrated strong performance in predicting protein structure, function, and other properties from individual sequences. To fully leverage ESM’s powerful encoding capabilities, we use it as the backbone of our protein encoder. Given a target protein sequence $\mathcal{P}$, we input it into the pre-trained ESM model and extract the final hidden state as the initial feature matrix:
@@
$\mathbf{X}_{p}=\text{AvgPool}_{s}(\mathbf{M}_{p}),$ (2)
where $\mathbf{X}_{p}\in\mathbb{R}^{D_{p}\times M}$, with $M=K/s$ denoting the number of local fragments. Next, a multi-layer deep neural network (DNN) is employed to adapt these features to our task and extract higher-level abstractions of local protein fragments. The encoding process is defined as:
@@
where $\mathbf{W}_{n}^{(l)}$ and $\mathbf{b}_{n}^{(l)}$ are the learnable weights and biases of the $l$-th DNN layer, $\mathbf{H}_{p}^{(l)}$ is the $l$-th hidden protein representation, and $\sigma(\cdot)$ is a non-linear activation function (ReLU in our case). The initial input is set as $\mathbf{H}_{p}^{(0)}=\mathbf{X}_{p}$. The final output, $\mathbf{H}_{p}\in\mathbb{R}^{D\times M}$, is adopted as the protein representation, where $D$ is the representation dimension.
@@
For each ligand compound, we convert its SMILES string into a 2D molecular graph $\mathcal{G}$. To represent the node (atom) information in $\mathcal{G}$, we initialize each atom using its chemical properties, following the implementation in the DGL-LifeSci library [53]. Each atom is initially represented as a 74-dimensional integer vector describing eight pieces of information: atom type, degree, number of implicit hydrogens, formal charge, number of radical electrons, hybridization state, number of total hydrogen atoms, and aromaticity. This initialization is formally written as:
@@
resulting in the initial node feature matrix $\mathbf{M}_{d}\in\mathbb{R}^{74\times N}$, where $N$ is the maximum allowed number of atoms per molecule. Molecules with fewer atoms are padded with virtual nodes filled with zeros to ensure consistent dimensions for batch training. Next, a linear transformation is applied to project the integer feature matrix into a real-valued one:
@@
producing a dense feature matrix $\mathbf{X}_{d}\in\mathbb{R}^{D_{d}\times N}$, with $D_{d}$ denoting the feature dimension.
@@
where $\mathbf{W}_{g}^{(l)}$ and $\mathbf{b}_{g}^{(l)}$ are the learnable weight matrix and bias vector of the $l$-th GCN layer, $\mathbf{H}_{d}^{(l)}$ is the $l$-th hidden node representation, $\mathbf{A}$ is the adjacency matrix of the graph $\mathcal{G}$ (augmented with self-loops), and $\sigma(\cdot)$ denotes a non-linear activation function (ReLU in our case). The input to the first layer is set as $\mathbf{H}_{d}^{(0)}=\mathbf{X}_{d}$. The final output, $\mathbf{H}_{d}\in\mathbb{R}^{D\times N}$, is adopted as the ligand representation.
@@
Thus far, the two encoders described above only capture the 1D and 2D structural information of proteins and ligands, respectively. To further enhance these encoders with biochemical factual knowledge, we integrate their outputs into a KG embedding framework. Specifically, we treat the encoded protein and ligand representations as entity embeddings within our constructed KG and incorporate them into a joint learning process alongside other entities and relations. The remainder of this subsection details our implementation.
Our KG structurally consists of entity-relation-entity triplets to describe relational knowledge. We define a triplet as $(h,r,t)$, where $h$ and $t$ are head and tail entities, and $r$ is the relation linking them, whose type is usually predefined by the schema. This KG is composed of two distinct subgraphs: one centered around proteins and the other around ligands.
  • Protein KG (GO-based): This subgraph incorporates two types of entities: protein entities $e_{protein}$ involved in the affinity prediction task, and Gene Ontology (GO) entities $e_{GO}$ describing functional annotations. The GO entities cover three categories: Molecular Function (MF), Cellular Component (CC), and Biological Process (BP). Therefore, the protein-related triplets fall into three groups: protein-MF, protein-CC, and protein-BP.
  • Ligand KG (LP-based): Similarly, this subgraph includes ligand entities $e_{ligand}$ and ligand property (LP) entities $e_{LP}$. The LP entities are further divided into Molecular Descriptors (MD), i.e., numerical attributes derived from molecular structures (e.g., counts of hydrogen bond donors/acceptors), and Chemical Features (CF), which denote qualitative attributes (e.g., hydrophobe, positively charged). Accordingly, the ligand-related triplets fall into two groups: ligand-MD and ligand-CF.
Next, we involve the protein and ligand representations in the KG embedding objective to enhance the protein and ligand encoding with external knowledge. Since the encoders yield matrix-form local representations unsuitable for KG embeddings, we first compute vector-form global representations by averaging across the local representations, then project them into the KG embedding space via a linear transformation:
@@
where $\text{Average}(\cdot)$ denotes the averaging operation over all local representations. $\mathbf{W}_{k}^{*}$ and $\mathbf{b}_{k}^{*}$ are learnable weight matrices and bias vectors that map the global representations into the KG embedding space. The resulting $\mathbf{h}_{p}$ and $\mathbf{h}_{d}\in\mathbb{R}^{D}$ serve as the semantic embeddings of the protein and ligand entities in the KG. Similar to [54], the embedding sets of other entities and relations are denoted by learnable $\mathbf{T}\in\mathbb{R}^{D\times T}$ and $\mathbf{R}\in\mathbb{R}^{D\times R}$, where $T$ is the number of non-protein and non-ligand entities, and $R$ is the number of relation types.
To achieve the KG embedding objective, we train the KG by optimizing the corresponding embeddings to fit the triplet relations. The objective function is defined as:
$\mathcal{L}_{KGE}=\frac{1}{|\mathcal{S}|}\sum_{(h,r,t)\in\mathcal{S}}F(h,r,t),$ (8)